OpenStudy (anonymous):

3+5+7+.....+k=440 how i find k?

OpenStudy (anonymous):

is about aritmetic?

OpenStudy (anonymous):

oops sorry

OpenStudy (anonymous):

we have an arithmetic Sequence with d = 2 , a1=3 the sum of n terms is : (n/2) * (2*a1 + (n-1)*d) = 440

OpenStudy (anonymous):

now when we will have 'n' we can find k since it's the nth element an = a1 + (n-1)*d = k

OpenStudy (anonymous):

are you there ?

OpenStudy (anonymous):

yes iam...

OpenStudy (anonymous):

so are you trying to solve it using what i said ?

OpenStudy (anonymous):

now...i wil try to find

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

where is n?

OpenStudy (anonymous):

you have to find it.. using (n/2) * (2*a1 + (n-1)*d) = 440

OpenStudy (anonymous):

n is the index of the number k you are looking for

OpenStudy (anonymous):

still understand with n

OpenStudy (anonymous):

what ?

OpenStudy (anonymous):

in my case ,value n is ?

OpenStudy (anonymous):

n is the index of the number k .. a1 = 3 , a2 = 5 ,... , an=k

OpenStudy (anonymous):

weneed to find n using : (n/2) * (2*a1 + (n-1)*d) = 440 then .. when we have n we can find k using : k = an = a1 + (n-1)*d

OpenStudy (anonymous):

ok still try

OpenStudy (anonymous):

2n+12(n-1)=440..then?

OpenStudy (anonymous):

n=328/12

OpenStudy (anonymous):

this is not correct.. (n/2) * (2*a1 + (n-1)*d) = 440 (n/2) * (2*3 + (n-1)*2) = 440 (n/2) * (4 + 2n) = 440 n * (2+ n) = 440 n^2 + 2n = 440

OpenStudy (anonymous):

this will give you two solution which one of them will be negative and thus not the right one since n is index it is positive

OpenStudy (anonymous):

so i choose positif index?

OpenStudy (anonymous):

yes.. what is n from this equation ?

OpenStudy (anonymous):

then enter to an=al-(n-1)d=440 ?

OpenStudy (anonymous):

if we found n=20...then a(20)=2-(20-1)3====)an=-55?

OpenStudy (anonymous):

an = a1 + (n-1)*d = k

OpenStudy (anonymous):

3 + (20 -1 )*2 = 41

OpenStudy (anonymous):

there is a plus not a minus between a1 and (n-1)d

OpenStudy (anonymous):

you should review the basics of arithmetic sequences ..

OpenStudy (anonymous):

ok .thanks brother....

OpenStudy (anonymous):

yw:)